nila:

The International Business Club starts the school year with $250.50 in their account. They spend $35 each month on activities. The Future Agricultural Leaders Club starts with $300 and spends $45.25 each month. Which equation can be used to find m, the number of months it will take for both accounts to have the same amount of money? 250.5 + 35m = 300 + 45.25m 250.5 – 35m = 300 – 45.25m 250.5m – 35 = 300m – 45.25 250.5m + 35 = 300m + 45.25

Bigbosssaint21:

I feel like it could be one of the first two

dude:

Ya Separate the values between clubs International Business Club (IBC) 35/mo 250.50 starting Future Agricultural Leaders Club (FALC) 45.25/mo 300 starting mo = months, so IBC = 250-.50-35m FALC = 300-45.25m These numbers should looks familiar in the choices